Fermi National Laboratory


Accelerator Update for March 18, 2005 - March 21, 2005

Friday March 18
The day shift began with Operations monitoring store 4046 and a small Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

Operations terminated store 4046 at 9 AM.

Operators prepared the Collider Detector at Fermilab's (CDF) collision hall for an access that began at 9:23 PM. The experimenters were out of access by 10:16 AM.

Operations put the Tevatron (TeV) into a dry squeeze at 10:40 AM.

Operations put the TeV into shot setup at 11:05 AM.

Operations established store 4048 at 1:20 PM with an initial luminosity of 113.143E30. This store set a New Luminosity Record. Fermilab averages the luminosities recorded by CDF and D-Zero (D0) to find the initial luminosity. CDF recorded a luminosity of 120.298E30 and D0 recorded 105.988E30. Both experiments broke previous bests with this store, but this was the first time D0 had ever recorded a luminosity above the 100E30 level.

Pbar resumed stacking at 1:28 PM.

MCenter resumed taking beam at 1:36 PM.

At 1:58 PM, Operations reported that the Meson MTest beam was being held off due to wild oscillations by a power supply (MT2WD1). Operators contacted an expert, but before he ever got the chance to look at the supply it stopped oscillating. Operators established beam to MTest at 2:10 PM.

At 3:36 PM, the TeV suffered a multi-house quench due to a kicker (AAK1) prefire.

CDF briefly accessed their collision hall at 4:57 PM, came out, and then accessed their hall again an hour later.

Kicker technicians began repairs on AAK1 at 5:39 PM.

MTest experimenters reported that they were finished with the beam for the weekend at 7:01 PM.

Operations put the Recycler into shot setup at 7:56 PM to stash antiprotons. Any shot setup halts stacking and beam to NuMI and Switchyard (SY).

The Recycler completed their stashing and Pbar resumed stacking at 9:13 PM. MCenter resumed taking beam a few minutes later.

At 10:23 PM, a Linac RF station (KRF5) tripped off and would not reset due to a drop in temperature. Operators investigated and then called in an expert.

Saturday March 19
The midnight shift began with Operations monitoring the stack and stash, but with stacking, MiniBooNE, NuMI, and, MCenter off due to KRF5 work.

The Linac expert manually set the temperature regulation and got KRF5 back on at 12:55 AM, but he contacted another expert for further help.

At 1:18 AM, a Booster RF station (BRF4) tripped off and wouldn't reset. Operators decided to give it a little rest.

Operations put the TeV through a dry squeeze at 1:58 AM.

At 2:43 AM, a Main Injector RF station (MIRF2) tripped off and wouldn't reset. Operators bypassed the station.

The kicker expert reported at 2:44 AM that the AAK1 kicker was on and pulsing normally. It will need to condition for eight hours. He put all the kickers into conditioning mode during this time period.

At 9:53 AM, MCenter reported beam problems. Operators discovered a power supply (MC5Q1) had tripped off. They replaced a control card and had it back on by 10:26 AM.

Operations reported at 10:04 AM that an expert had MIRF2 back on line. He also said that this station would need more work soon.

At 10:19 AM, Operations reported that the Linac drift tube tank one tripped off due to beam problems for the first of four times in five minutes. It reset okay each time.

Operations ramped the TeV to 150 GeV at 11:11 AM. At the same time a TeV sector B4 wet engine came into alarm. Operators asked cryo system techs to investigate and make repairs.

Operations set the AAK1 kicker to its normal voltage at 12:10 PM.

At 12:33 PM, cryo techs reported that the B4 wet engine repair would take about two hours.

At 12:47 PM, AAK1 tripped off for the first of many times over the next five hours.

A TeV expert began a study period at 12:53 PM.

At 12:59 PM, MC5Q1 tripped off again. Operators couldn't reset the supply and called in an expert to make repairs.

Operations reported at 4:14 PM that the TeV B4 work was finished and that they were beginning to ramp the TeV to 980 GeV.

Operations established beam to MCenter at 4:29 PM.

TeV experts began another study period at 5:29 PM.

At 6:24 PM, Operations reported that the TeV quenched due to the study.

Experts completed a full Booster tune at 7:41 PM.

Operation reported at 7:47 PM that operators began the first of three TeV QBS ramps.

At 8:02 PM, the TeV abort kickers fired during a dry squeeze test. They reset okay.

TeV experts had another study period from 8:58 PM until about 10:34 PM.

Operations halted MCenter beam 11:07 PM in preparation of shot setup.

Sunday March 20
The midnight shift began with Operations monitoring the stack and stash, with MiniBooNE taking beam, but with MCenter off due to shot setup.

At 1:12 AM, Operations reported that shot setup was delayed because of a Pbar sequencer problem. Operators investigated and found that a scope used to check emittances was missing.

At 1:24 AM, KRF5 began tripping off and continued to do so for the next twenty minutes.

Operations established store 4052 at 3:10 AM with an initial luminosity of 104.991E30. Pbar resumed stacking.

NuMI and MCenter resumed taking beam at 3:37 AM.

Operations put the Recycler into shot setup at 7:19 AM and stashing began about fifteen minutes later.

Operators completed the Recycler stash at 9 AM and resumed sending beam to Pbar, NuMI, and MCenter.

At 10:02 AM, Operations reported that MIRF2 tripped off for the first of a number of times over the next three hours, and that MIRF9, 17, & 18 took turns joining in with those trips.

A Pbar expert began Debuncher cooling study at 6:12 PM that would slow stacking.

At 9:10 PM, Operations reported that the AP50 Pbar vacuum readbacks disappeared, which caused the beam permit to trip off. Operator investigated and then called experts for help.

Operations switched the MI NuMI/Pbar slipstacking to a NuMI only mode at 9:51 PM.

At 10:28 PM, Operations reported losing the NuMI permit due to a bad input to the Absorber's cooling system. Operator called an expert for advice.

Controls experts reported at 11:24 PM that a crate power supply for AP50 vacuum crate had failed. They were beginning to replace the supply.

Operators made up the water for the NuMI absorber at 11:25 PM.

Pbar resumed stacking around 11:45 PM.

Monday March 21
The midnight shift began with Operations monitoring store 4052 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 12:06 AM, a RF problem caused MI to loose its beam permit. Operators had the permit reset by 12:13 AM.

At 5:06 AM, NuMI began tripping off due to beam problems. Operators resumed the NuMI/Pbar slipstacking mode at 5:38 AM, which resolved the problem.

The Plan for Monday
The plan for today is to terminate store 4052 and then establish a new store around 10:30 AM. The rest of the day will include stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/21/2005   email Fermilab